A mobile game is such that the character will always walk 3 steps North, 4 steps East, 11 steps

South and 10 steps West in the same order, over and over again. How far is the character from

its original position after doing the procedure 20 times


1
Expert's answer
2021-01-28T05:24:19-0500

After the first procedure character is 11-3=8 steps South and 10-4=6 steps West.

After 20th procedure character is 8*20=160 steps South and 6*20=120 steps West.

Distance from the origin: "\sqrt{160^2+120^2}=\sqrt{40000}=200" steps.